A proton accelerates from rest in a uniform electric field of 5NC^-1. At some time later it's speed is 1.2×10^6 ms^-1. What is the acceleration of the proton


1
Expert's answer
2022-05-09T14:49:43-0400
"a=\frac{F}{m}=\frac{qE}{m}"

"a=\frac{1.6*10^{-19}*5}{9.1*10^{-31}}=8.8*10^{11}\:\rm m/s^2"